OpenStudy (anonymous):

Can someone help me out with functions? How do you solve a function set up as f(g(x))? I need help with the concept, not a specific problem.

OpenStudy (anonymous):

specific problem to help you explain, though: Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

random231 (random231):

okay : now f(x)=9x-2 this is a function of x (that is why x is in brackets) now when its f(g(x)): the function is of g(x) so basically you hav to substitute g(x) in place of x. according to your question: f(g(x))= 9(-x+3)-2

OpenStudy (anonymous):

Oh, okay...so this would be...-9x + 25?

OpenStudy (anonymous):

It's been over a year since I've had a math class, and I had a handle on these a few weeks ago, but I had to drop and re-enroll in the course (algebra II), so I lost all my new understanding of them. Lol. That just brought it back, though.

random231 (random231):

yes u r right! sorry i lost net connection!

OpenStudy (anonymous):

Thanks!! You helped a lot!

OpenStudy (anonymous):

Could you help with another brand of functions?

random231 (random231):

sure!

OpenStudy (anonymous):

How do you solve something where f is raised to the power of -1?

OpenStudy (anonymous):

f^-1(x) is an inverse function, right?

random231 (random231):

yup!

OpenStudy (anonymous):

How do you solve that? e.g. Let f(x) = x2 - 16. Find f-1(x).

OpenStudy (anonymous):

x^2 and f^-1x, btw.

random231 (random231):

okay:

random231 (random231):

ill give u another example and will solve this one!

OpenStudy (anonymous):

Thanks so much :)

random231 (random231):